OpenStudy (anonymous):

How much interest is earned for the investment? $20,000 for 2 years at 6% compounded annually

OpenStudy (jdoe0001):

do the same as before \(\bf Balance=P\left(1+\frac{r}{n}\right)^{nt}\\ \quad \\ P=principal\\ r=rate\ 6\% = 0.06\\ n=period\textit{per year, 12 months per year, }=12\)

OpenStudy (mathstudent55):

Calculate the future value, then subtract the present value from it. The difference is the interest.

OpenStudy (anonymous):

i got 224.72

OpenStudy (jdoe0001):

ohh ahemm. right... tis the interest only =)

OpenStudy (mathstudent55):

F = P(1 + r)^t F = 20,000(1 + 0.06)^2 F = 22,472 22,472 - 20,000 = 2472

OpenStudy (anonymous):

did i get it righjt

OpenStudy (mathstudent55):

Are you sure you used 20,000 in your calculation? It seems you entered 2000 instead.

OpenStudy (mathstudent55):

You're off by a factor of 10 The answer is 2472 You have 247.2

OpenStudy (anonymous):

ya tht wht i got